# Cognitive Paradigms for Evaluating VLMs on Visual Reasoning Task

### v1

@@ Abstract
Evaluating the reasoning capabilities of Vision-Language Models (VLMs) in complex visual tasks provides valuable insights into their potential and limitations. In this work, we assess the performance of VLMs on the challenging Bongard Openworld Problems benchmark, which involves reasoning over natural images. We propose and evaluate three human-inspired paradigms: holistic analysis (global context processing), deductive rule learning (explicit rule derivation and application), and componential analysis (structured decomposition of images into components). Our results demonstrate that state-of-the-art models, including GPT-4o and Gemini, not only surpass human benchmarks but also excel in structured reasoning tasks, with componential analysis proving especially effective. However, ablation studies reveal key challenges, such as handling synthetic images, making fine-grained distinctions, and interpreting nuanced contextual information. These insights underscore the need for further advancements in model robustness and generalization, while highlighting the transformative potential of structured reasoning approaches in enhancing VLM capabilities.
@@ 1 Introduction
Human intelligence seamlessly integrates visual perception with language and higher-level cognition, enabling reasoning about visual information Kunda (2020). This ability, known as visual reasoning, supports tasks such as understanding spatial relationships, inferring object functions, and grounding language in the physical world. Despite advancements in natural language processing (NLP) Zhao et al. (2023), achieving robust visual reasoning in machines remains a significant challenge, particularly in multimodal contexts Cao et al. (2024).
@@
Large Language Models (LLMs) have transformed NLP by excelling in tasks like machine translation, text generation, and question answering Brown et al. (2020); Devlin (2018). Their emergent abilities in areas such as code generation Chen et al. (2021) and logical reasoning Wei et al. (2022) suggest potential for bridging language and vision to tackle complex visual reasoning tasks.
@@
Visual reasoning tasks range from basic object recognition Lin et al. (2014) to abstract inference Zerroug et al. (2022); Teney et al. (2020); Barrett et al. (2018); Zhang et al. (2019). Complex tasks demand integrating perceptual data with higher-order cognition, such as spatial reasoning and relational understanding Cao et al. (2024); Vaishnav et al. (2022). Bongard Problems (BPs) Bongard (1968), involving abstract rules distinguishing two image sets, exemplify such challenges by requiring both perceptual and conceptual reasoning.
Structured reasoning tasks like BPs require the reasoning process to follow specific rules or structures. These tasks involve identifying patterns or relationships among image components and applying abstract reasoning based on those patterns. Vision-Language Models (VLMs), such as Phi-1.5 Javaheripi et al. (2023), Llama-vision Meta (2024), Gemini Google (2024), and GPT-4o OpenAI (2024), provide a robust foundation for tackling BPs. By leveraging multimodal datasets, these models combine visual perception with higher-order cognitive processes, making them well-suited for structured reasoning tasks like BPs. This capability allows them to approach complex visual reasoning by breaking down information into identifiable, logical steps.
Figure 1: Illustration of a Bongard OpenWorld dataset showing three positive and negative samples along with a query image. Distinguishing rule in this example is A group photo at a wedding reception.. Based on the rule, query image belongs to negative category.
This paper investigates the ability of Vision-Language Models (VLMs) to solve natural image-based Bongard Problems using the Bongard Openworld Dataset Wu et al. (2024) (example shown in Figure 1). Inspired by human problem-solving strategies, we evaluate three paradigms: holistic analysis, deductive rule learning, and componential analysis. Our findings demonstrate that VLMs employing these strategies can surpass human performance on this challenging visual reasoning task. Additionally, through ablation studies and a comparative analysis of open- and closed-source models, we identify key limitations in multi-image reasoning, such as difficulties with fine-grained distinctions, synthetic images, and nuanced contextual understanding. These insights provide critical directions for advancing the robustness and human-like reasoning capabilities of VLMs.
@@
    We evaluate the effectiveness of Vision-Language Models (VLMs) on the challenging task of solving Bongard Openworld Dataset—natural image-based BPs.
@@
    We propose and evaluate three distinct paradigms for applying VLMs to this task, inspired by human learning and problem-solving strategies:
    1. 2.1.
      Holistic Analysis (Analyzing all information together)
    2. $2.$
      Deductive Rule Learning (Focusing on deriving rules before testing scenarios)
    3. 2.3.
      Componential Analysis (Decomposing images into structured, interpretable components)
@@
    We demonstrate that VLMs can surpass human performance in this challenging open-world visual reasoning task using human-centric learning methodologies.
@@
    We provide a detailed evaluation of both open-source and closed-source models, highlighting current limitations and performance discrepancies in structured reasoning tasks.
  5. 5.
    We demonstrate that addressing limitations in fine-grained distinctions, synthetic image processing, and contextual understanding is key to improving VLM robustness and aligning them more closely with human reasoning capabilities.
@@ 2 Related Work
To evaluate the capabilities of VLMs, various benchmarks have been developed, focusing on different aspects of visual understanding and reasoning. These benchmarks play a crucial role in assessing model performance and identifying areas for improvement. Traditional benchmarks focussed on single image tasks like image captioning Lin et al. (2014), OCR Mishra et al. (2019); Singh et al. (2021), and VQA Antol et al. (2015); Zhang et al. (2024); Lu et al. (2022); Gurari et al. (2018). For multi-image comprehension, researchers have developed interleaved image-text corpora, including MMC4 Zhu et al. (2024) and OBELICS Laurençon et al. (2024); Lin et al. (2024), along with dedicated benchmarks like Q-Bench Wu et al. (2023), Memontos Wang et al. (2024), and DEMON Li et al. (2023). While these benchmarks have contributed to progress in evaluating multi-image understanding, they often lack comprehensive coverage of complex reasoning scenarios. Other benchmarks like MMIU (Multimodal Multi-Image Understanding) Meng et al. (2024) have been pivotal in evaluating VLMs comprehensively, incorporating diverse datasets covering areas like science and mathematics.
@@
Despite advancements in VLM benchmarks that involve language, purely visual reasoning tasks without linguistic input remain crucial for evaluating core visual understanding. Bongard problems Bongard (1968) present a unique challenge in this regard. Their significance lies in their focus on abstract visual concepts, the absence of explicit language, and the inherent multi-image relational reasoning required for their solution. Extensions such as Bongard-HOI Jiang et al. (2022) and Bongard-OpenWorld Wu et al. (2024), further extend these challenges to human-object interactions and real-world scenarios, highlighting models’ few-shot reasoning capabilities.
@@
Beyond Bongard problems, other purely visual reasoning challenges have been proposed, each focusing on different aspects of visual cognition. These include Synthetic Visual Reasoning Test (SVRT) Fleuret et al. (2011), Compositional Visual Reasoning (CVR) Zerroug et al. (2022), Raven’s Progressive Matrices (RPM) Barrett et al. (2018); Zhang et al. (2019), Visual Progressive Matrices (V-PROM) Teney et al. (2020), and the Abstract Reasoning Corpus (ARC) Chollet (2019). These benchmarks collectively assess a range of visual reasoning abilities, including spatial reasoning, compositional reasoning, and abstract pattern recognition, providing valuable insights into the visual understanding capabilities of AI models.
Human cognitive processes offer valuable insights for evaluating visual reasoning in VLMs. Three primary learning strategies are particularly relevant: holistic analysis, deductive analysis, and componential analysis. Holistic analysis involves integrating all available information simultaneously, akin to similarity-based strategies where new information is compared to existing knowledge von Helversen et al. (2014); Johnson-Laird (2010). This approach emphasizes understanding the overall context and interconnections between elements. Focused problem-solving prioritizes understanding underlying principles through deductive reasoning before application, mirroring rule-based strategies Knauff and Wolf (2010); Helie et al. (2015); Rips (1994). This strategy emphasizes abstract rule extraction and application.
Finally, componential analysis breaks down problems into individual components, extracting relevant details before integration Wu and Molnár (2022); Gluck et al. (2008); Hutchinson (1992). This approach emphasizes detailed analysis of individual elements and their subsequent combination. These distinct human learning strategies provide a framework for designing and interpreting visual reasoning benchmarks for VLMs, allowing researchers to assess how well these models emulate different aspects of human cognition.
@@ 3 Models
To evaluate our approach comprehensively, we employed a diverse range of Vision-Language Models (VLMs) that include both closed-source and open-source architectures. This diversity allowed us to assess the strengths and limitations of different model types and scales.
This selection encompasses state-of-the-art multimodal models, including GPT-4o OpenAI (2024), Gemini-2.0 Google (2024), Pixtral-12B Agrawal et al. (2024), Llama-Vision-3.2 (11B, and 90B) Meta (2024), LLaVA (7B, 13B and 34B) Liu et al. (2023), and LLaVA built on Llama 3 (8B) XTuner (2025). By leveraging this variety, we systematically analyzed model performance across different architectural designs and parameter scales, offering a holistic view of their capabilities. For the evaluation purpose, two metrics are used: classification accuracy and semantic similarity (more in Appendix C).
Further technical details, including model architectures and individual summaries, are elaborated in Appendix D.
@@ 4 Dataset
We evaluate our approach on the Bongard OpenWorld dataset Wu et al. (2024), a benchmark designed to assess few-shot visual reasoning abilities. The full dataset contains 1001 cases, from which we have chosen a sample of 500 cases for our experiments. Each case in the dataset consists of six positive and six negative examples and a query image. These cases are governed by an underlying "commonsense" rule, with the rules themselves categorized into distinct visual concepts (further details can be found in Appendix B). Figure 1 illustrates the task: given a set of positive and negative examples, the goal is to determine the underlying rule and classify a query image accordingly.
@@ 5 Experiments
This paper investigates VLMs’ ability to solve natural image-based Bongard Problems. We evaluate three paradigms: holistic analysis (processing all information together), deductive rule learning (deriving rules first), and componential analysis (examining components separately).
5.1 Holistic Analysis
In this section, we describe our experiments using a holistic analysis approach, where the model is presented with all images (positive, negative, and test) simultaneously.
The holistic analysis experiments assess the ability of VLMs to directly infer patterns and rules from a set of images. We designed a prompt that presents the model with m positive examples, n negative examples, and a final query image.
The prompt instructs the model to:
  1. 1.
    Identify a rule that distinguishes the positive examples from the negative examples;
  2. 2.
    Analyze the query image;
  3. 3.
    Classify the query image as either positive or negative based on the identified rule.
It is explicitly requested to the VLM model to provide its analysis, the derived rule, a description of the query image, and the final classification category. Details are provided in Appendix E.1.
@@ 5.2 Deductive Rule Learning
This section describes our experiments using a deductive rule learning approach, which involves a two-step process: first, the model infers a general rule; second, it applies this rule to classify a query image.
The deductive rule learning experiments assess the ability of VLMs to explicitly formulate a rule before classifying a query image. This approach mirrors human deductive reasoning. We employed a two-stage prompting strategy. In the first stage, we presented the model with m positive examples and n negative examples and asked it to identify a rule that distinguishes the two categories. The prompt requested a concise summary of the identified rule (less than 20 words). The precise wording of this first-stage prompt is provided in Appendix E.2.
In the second stage, we presented the model with the previously generated rule summary and a query image. The prompt instructed the model to analyze the query image in the context of the given rule and classify it as either positive or negative. The prompt also requested a brief analysis of the query image and a restatement of the rule. The precise wording of this second-stage prompt is provided in Appendix E.2.
@@ 5.3 Componential Analysis
This section describes our experiments using a componential analysis approach (detailed in Appendix E.3). This method involves decomposing each image into its constituent visual elements and then reasoning based on these components (textual elements). This process is divided into three main stages: (1) detailed image description generation, (2) providing instruction on rule derivation, and (3) rule extraction and application based on these descriptions.
The componential analysis experiments aim to evaluate the ability of VLMs to reason based on structured descriptions of visual content. This approach simulates how humans might break down complex visual scenes into simpler parts for analysis. We employed a three-stage prompting strategy.
In the first stage, we generate detailed descriptions of each image. This prompt instructs the model to analyze an image and provide a structured description, covering aspects such as scene, objects (living and inanimate), activities, contextual elements, visual patterns, emotional undertones, textual information, and a summary. In the second stage, we provide the models with instructions on how to analyze the outputs and derive the rule.
In the final stage, which is purely textual, we provided the model with the descriptions generated in the first stage, along with information about which images belong to positive examples and negative examples, and the query image. The prompt then asks the model to derive a rule based on the descriptions of positive and apply this rule to classify the query image.
5.4 Rule-Based Evaluation
This section describes experiments designed to disentangle the image analysis (perception) and rule application (abstract reasoning) capabilities of VLMs. Unlike the previous experiments, where the models were tasked with both extracting and applying rules, here we provide the models with pre-extracted rules and assess their ability to apply these rules to classify query images.
These experiments isolate the rule application component of the Bongard problem-solving process. This experimental setup is conceptually similar to the Deductive Rule Learning experiments (Section 5.2), which also involved a two-step process: (1) inferring a general rule and (2) applying that rule to classify a query image. However, in this experiment, we bypass the first step (rule inference) and instead use the rule summaries generated by GPT-4o in the Deductive Rule Learning experiment. For each Bongard problem, we took the rule summary generated by GPT-4o. We then presented the model with the query image and the corresponding rule summary, explicitly instructing it to classify the query image based on the provided rule. This approach allows us to specifically assess how well the models can understand and apply a given rule, independent of their ability to extract the rule themselves.

### v2

@@ Abstract
Advancing machine visual reasoning requires a deeper understanding of how Vision-Language Models (VLMs) process and interpret complex visual patterns. This work introduces a novel, cognitively-inspired evaluation framework to systematically analyze VLM reasoning on natural image-based Bongard Problems. We propose three structured paradigms—Direct Visual Rule Learning, Deductive Rule Learning, and Componential Analysis—designed to progressively enforce step-wise reasoning and disentangle the interplay between perception and reasoning. Our evaluation shows that advanced, closed-source VLMs (GPT-4o, Gemini 2.0) achieve near-superhuman performance, particularly when provided with high-quality image descriptions, while open-source models exhibit a significant performance bottleneck due to deficiencies in perception. An ablation study further confirms that perception, rather than reasoning, is the primary limiting factor, as open-source models apply extracted rules effectively when given accurate descriptions. These findings underscore the critical role of robust multimodal perception in enhancing generalizable visual reasoning and highlight the importance of structured, step-wise reasoning paradigms for advancing machine intelligence.
@@ 1 Introduction
Human intelligence seamlessly integrates visual perception with language and higher-level cognition, enabling sophisticated reasoning about the visual world Kunda (2020). This ability, known as visual reasoning, supports tasks ranging from understanding spatial relationships to inferring object functions. Despite advancements in natural language processing (NLP) Zhao et al. (2023), achieving robust visual reasoning in machines, especially in complex, multimodal contexts, remains a significant challenge Cao et al. (2024).
@@
Large Language Models (LLMs) have revolutionized NLP, demonstrating impressive capabilities in tasks like machine translation, text generation, and question answering Brown et al. (2020); Devlin (2018). Their emergent abilities in areas such as code generation Chen et al. (2021) and logical reasoning Wei et al. (2022a) suggest their potential for tackling complex visual reasoning tasks. However, effectively bridging language and vision to achieve human-like visual reasoning requires careful consideration of how these models process and interpret visual information.
@@
Visual reasoning challenges span a spectrum from basic object recognition Lin et al. (2014) to abstract inference Barrett et al. (2018); Zhang et al. (2019); Teney et al. (2020); Zerroug et al. (2022). One such challenge – Bongard Problems (BPs) Bongard (1968), involving abstract rules distinguishing two image sets, exemplify the challenges of integrating perceptual data with higher-order cognition. They require both perceptual and conceptual reasoning, making them a valuable testbed for evaluating visual reasoning capabilities.
Vision-Language Models (VLMs), such as Phi-1.5 Javaheripi et al. (2023), Llama-vision Meta (2024), Gemini 2.0 Google (2024), and GPT-4o OpenAI (2024), offer a promising avenue for addressing BPs. By leveraging multimodal training data, these models combine visual perception with higher-order cognitive processes. However, the mechanisms by which VLMs achieve visual reasoning, and how these mechanisms align with human cognitive strategies, remain an open question.
This paper investigates the visual reasoning abilities of VLMs on natural image-based Bongard Problems using the Bongard Openworld Dataset Wu et al. (2024). Inspired by human problem-solving strategies, we introduce three distinct paradigms—direct visual rule learning, deductive rule learning, and componential analysis—designed to dissect the interplay between perception and reasoning in VLMs. Our key contribution is the development and application of these cognitively-inspired paradigms, which allow us to isolate and analyze the distinct roles of perception and reasoning in VLM performance. We demonstrate that advanced VLMs can achieve near- and even super-human performance on this challenging task, but also reveal critical limitations in the perception modules of open-source VLMs. These findings offer valuable insights into the computational underpinnings of visual reasoning and provide crucial directions for future research.
@@
    Three distinct, cognitively-inspired paradigms for evaluating VLM’s visual reasoning ability: Direct Visual Rule Learning, Deductive Rule Learning, and Componential Analysis, isolating the roles of perception and reasoning.
@@
    Demonstration of near- and super-human performance by VLMs, particularly with high-quality image descriptions.
@@
    Identification of a critical bottleneck in open-source VLMs: the perception module (image description generation), which significantly limits their reasoning capabilities.
@@
    Highlighting the crucial role of robust image descriptions for complex visual reasoning and the need for future research to focus on improving the perception components of open-source VLMs.
@@ 2 Related Work
Evaluating Vision-Language Models (VLMs) requires diverse benchmarks targeting various aspects of visual understanding and reasoning. Traditional benchmarks focused on single-image tasks like image captioning Lin et al. (2014), OCR Mishra et al. (2019); Singh et al. (2021), and VQA Antol et al. (2015); Zhang et al. (2024); Lu et al. (2022); Gurari et al. (2018). However, the field has increasingly shifted towards more complex scenarios, including multi-image comprehension. This shift has led to the development of interleaved image-text corpora like MMC4 Zhu et al. (2024) and OBELICS Laurençon et al. (2024); Lin et al. (2024), along with dedicated benchmarks specifically designed for multi-image understanding, such as Q-Bench Wu et al. (2023), Memontos Wang et al. (2024), and DEMON Li et al. (2023). Benchmarks like MMIU Meng et al. (2024) further broaden the scope by incorporating diverse datasets. While these benchmarks represent significant progress, they often fall short in capturing the intricacies of abstract visual reasoning, particularly those devoid of linguistic input.
@@
Purely visual reasoning tasks, where language is absent, are crucial for isolating and evaluating core visual understanding. Bongard problems Bongard (1968) provide an excellent example of such tasks, focusing on abstract visual concepts and requiring multi-image relational reasoning. Datasets like Bongard-HOI Jiang et al. (2022) and Bongard-OpenWorld Wu et al. (2024) extend these challenges to human-object interactions and real-world scenarios, respectively, pushing models to demonstrate robust few-shot reasoning capabilities.
@@
This focus on purely visual reasoning is not unique to Bongard problems. Other notable benchmarks in this category include SVRT Fleuret et al. (2011), CVR Zerroug et al. (2022), Raven’s Progressive Matrices (RPM) Barrett et al. (2018); Zhang et al. (2019), V-PROM Teney et al. (2020), and ARC Chollet (2019). These benchmarks collectively assess a wide range of visual reasoning abilities, including spatial, compositional, and abstract reasoning, offering valuable insights into the visual cognition of AI models.
Our work is also informed by research in human cognition, which has identified key learning strategies relevant to visual reasoning. We consider three key learning strategies that are relevant to visual reasoning: direct visual rule learning – integrating all information simultaneously, akin to similarity-based strategies von Helversen et al. (2014); Johnson-Laird (2010), deductive reasoning – prioritizing rule extraction and application Knauff and Wolf (2010); Helie et al. (2015); Rips (1994), and componential analysis – breaking down problems into individual components Wu and Molnár (2022); Gluck et al. (2008); Hutchinson (1992). These strategies inform the design and interpretation of our visual reasoning benchmarks, enabling us to assess how well VLMs emulate different facets of human cognition.
Finally, our work has connections to the growing body of research on Chain-of-Thought (CoT) prompting Kojima et al. (2022); Wei et al. (2022b); Huang and Chang (2023). While our focus is on visual reasoning and human cognitive strategies, CoT prompting, which encourages models to generate intermediate reasoning steps, shares a similar goal of improving model reasoning abilities. Recent work has explored applying CoT to multimodal tasks Zhang et al. (2023); Chen et al. (2024); Shao et al. (2024), demonstrating its potential for enhancing performance. Our work, by explicitly separating to visual modality alone focusing on perception and reasoning and analyzing the models’ ability to extract and apply rules, provides a complementary perspective on understanding and improving visual reasoning in VLMs.
@@ 3 Models
To evaluate our approach comprehensively, we employed a diverse range of Vision-Language Models (VLMs) that include both closed-source and open-source architectures. This diversity allowed us to assess the strengths and limitations of different model types and scales. Our selection encompasses state-of-the-art multimodal models, including GPT-4o OpenAI (2024), Gemini 2.0-flash-exp, henceforth Gemini 2.0 Google (2024), Pixtral-12B Agrawal et al. (2024), Llama-Vision-3.2 (11B, and 90B) Meta (2024), LLaVA (7B, 13B and 34B) Liu et al. (2023), and LLaVA built on Llama 3 (8B) XTuner (2025). By leveraging this variety, we systematically analyzed model performance across different architectural designs and parameter scales, offering a holistic view.
For the evaluation purpose, two metrics are used: classification accuracy and semantic similarity. Further technical details, including model architectures, evaluation and individual summaries, are elaborated in Appendix A.4.
@@ 4 Dataset
We evaluate our approach on the Bongard OpenWorld dataset Wu et al. (2024), a benchmark designed to assess few-shot visual reasoning abilities. The full dataset contains 1001 cases, from which we have chosen a sample of equally balanced 500 test samples for our experiments (further details can be found in Appendix A.3). Each case consists of six positive and six negative examples and a query image. These cases are governed by an underlying “commonsense” rule, with the rules themselves categorized into distinct visual concepts. Figure A.1 illustrates an example task where the identifying rule is to detect a group wedding photo.
@@ 5 Experiments
This paper investigates VLMs’ ability to solve natural image-based Bongard Problems. We evaluate three paradigms: direct visual rule learning (processing all information together), deductive rule learning (deriving rules first), and componential analysis (examining components separately). For each paradigm, the VLMs were prompted to provide four distinct outputs: (1) an overall analysis of the visual features and their relationships, (2) the derived rule that distinguishes the positive and negative categories, (3) a description of the query image, and (4) the final classification category for the query image. This consistent, structured output format facilitated a detailed examination of the models’ reasoning processes across all paradigms.
Figure 1: A Multi-Stage Evaluation Framework for Visual Reasoning. Direct Visual Rule Learning: Positive and negative examples, along with a query image, are processed concurrently to identify the distinguishing rule and categorize. Deductive Rule Learning: A rule is extracted from the positive and negative examples and then applied to classify the query image. Componential Analysis: Textual descriptions of the images are generated, and rule extraction and classification are performed based solely on these textual representations.
5.1 Direct Visual Rule Learning
We begin by exploring a holistic approach to visual reasoning. In this paradigm, the Vision-Language Model (VLM) is presented with all images relevant to a given Bongard problem – the positive examples, the negative examples, and the query image – simultaneously. This approach aims to mimic human problem-solving strategies where all available information is considered concurrently. The prompt instructs the model to: (1) provide the overall analysis, (2) identify a rule that distinguishes the positive examples from the negative examples; (3) analyze the query image; (4) classify the query image as either positive or negative based on the identified rule The prompt is detailed in Appendix A.5.1.
@@ 5.2 Deductive Rule Learning
A core aspect of human intelligence is the ability to deduce general rules from specific examples and then apply those rules to new situations. This paradigm emulates human deductive reasoning, involving two key steps: first, inferring a general rule from a set of positive and negative examples; second, applying this derived rule to classify a query image.
To investigate this capability in VLMs, we designed a two-stage prompting strategy. In the first stage, the model was presented with six positive examples and negative examples. The prompt, detailed in Appendix A.5.2.1, instructed the model to identify a distinguishing rule that separates the two categories. Critically, the model was asked to provide a concise summary of this identified rule, limited to a maximum of 20 words. This constraint encouraged the model to focus on the most salient features distinguishing the categories.
In the second stage, the model was provided with the previously generated rule summary and a query image. The prompt, detailed in Appendix A.5.2.2, instructed the model to analyze the query image in the context of the given rule and classify it as either positive or negative. The model was also asked to provide a brief analysis of the query image and restate the rule it was applying. This allowed us to examine not only the classification decision but also the model’s reasoning process.
@@ 5.3 Componential Analysis
While deductive rule learning operates at the level of abstract rules, componential analysis takes a more granular approach. This method decomposes each image into its constituent visual elements and reasons based on these components, represented textually. This approach mirrors how humans often break down complex visual scenes into simpler parts for analysis. The process unfolds in two distinct stages: (1) detailed image description generation and (2) rule derivation and application.
The componential analysis experiments aim to evaluate the ability of VLMs to reason based on structured, textual descriptions of visual content. We employed a two-stage prompting strategy to simulate this compositional reasoning process. In the first stage, detailed descriptions of each image (positive, negative, and query) were generated using the prompt detailed in Appendix A.5.3.1. This prompt instructs the model to analyze an image and provide a structured textual description covering aspects such as scene, objects (living and inanimate), activities, contextual elements, visual patterns, emotional undertones, textual information, and a summary.
In the second stage, the model was provided with the descriptions generated in the first stage, along with explicit “textual” information about which images correspond to positive examples and which correspond to negative examples, as well as the description of the query image. The prompt, detailed in Appendix A.5.3.2, then instructed the model to derive a rule based on the descriptions and apply this rule to classify the query image – described in textual form. This stage combines the processes of rule derivation and application, relying solely on the textual representations of the visual content. This isolates the models’ ability to reason about visual concepts based on their textual representations.